What was asked the derivation and how to do solve .Any derivation solve the trick

Grade:12

1 Answers

Parnav Bansal
5 years ago
Derivation is proof or solution of any formul÷a .
Let take an example =as we know that v=u+at
Proof or derivation = a =differentiation of velocity w.r.t time
a=dc/dc
Then DV=ADT
Integrating both side 
v-u=at
v=u+at
Hence proofed